OpenStudy (anonymous):

Solve the equation for f: e+f+g/5=h

OpenStudy (anonymous):

h-e-g/5

mathslover (mathslover):

\[\huge{f = h-\frac{g}{5}-e}\]

mathslover (mathslover):

\[\huge{f=h-e-\frac{g}{5}}\]

mathslover (mathslover):

\[\huge\color{violetred}{f=h−e−\frac{g}{5}}\]